### v1

@@ Abstract
Parameter calibration is crucial for reducing uncertainty and improving simulation accuracy in physics-based models, yet computational constraints pose significant challenges. Bayesian calibration methods offer a principled framework for combining prior knowledge with data while rigorously quantifying uncertainty. In this work, we compare four emulator-based Bayesian calibration methods—Calibrate-Emulate-Sample (CES), History Matching (HM), Bayesian Optimal Experimental Design (BOED), and a novel Goal-Oriented BOED (GBOED) approach—using the Lorenz ’96 multiscale system as a testbed. Our GBOED formulation explicitly targets calibration-relevant quantities and leverages information-theoretic criteria for data selection. We assess each method in terms of calibration accuracy, uncertainty quantification, computational cost, and convergence behavior. We evaluate each method’s performance in balancing computational cost, implementation complexity, and uncertainty quantification (UQ), with additional insights into convergence behavior as model evaluations increase. We find CES offers excellent performance but at high computational expense, while GBOED achieves comparable accuracy using fewer model evaluations. Standard BOED underperforms with respect to calibration accuracy, and HM shows moderate effectiveness but can be useful as a precursor. Our results highlight trade-offs among Bayesian strategies and demonstrate the promise of goal-oriented design in calibration workflows.
@@ 1 Introduction
Uncertainty quantification (UQ) plays a crucial role in validating computational models, particularly in climate science and other physics-based systems where predictions influence critical decision-making. Many-query problems, such as the calibration of expensive physics models, require repeated evaluations of complex simulations, making UQ both essential and computationally challenging. In these settings, calibration involves estimating uncertain model parameters, propagating their uncertainty through the system, and assessing the reliability of predictions. To be practically useful, UQ methods for calibration must not only be accurate but also computationally efficient and scalable.
A key approach to mitigating computational costs is the use of emulators, such as Gaussian processes (GPs), which approximate expensive simulations while retaining predictive accuracy and providing robust UQ. Emulators enable faster inference in many-query problems like calibration, but their effectiveness depends on the choice of training data. Selecting informative simulation runs to build the emulator is critical, as inefficient sampling can lead to poor calibration performance and unnecessary computational expense.
This paper examines Bayesian calibration of expensive models using emulators, focusing on different data acquisition strategies. We compare four key approaches: Bayesian Optimal Experimental Design (BOED) [33, 5, 20] and its goal-oriented extension (GBOED) [60, 1, 2]), Calibrate Emulate Sample (CES) [7], and History Matching (HM) [9]. Among these, optimal experimental design (OED) methods like BOED have been widely applied to guide physical experiments and sensor placement [58, 48, 18], but their potential for selecting computer simulations in emulator-based calibration remains underexplored.
Our contributions are twofold. First, we introduce a novel goal-oriented BOED (GBOED) algorithm that extends standard BOED to optimize information gain about the posterior distribution, making it more effective for calibration tasks. Second, we provide a comprehensive comparative study of BOED, GBOED, CES, and HM, evaluating their trade-offs in computational cost, implementation complexity, and UQ performance. Our results offer new insights into the practical merits and limitations of BOED-based methods for calibration and contribute to the broader discussion of how best to integrate experimental design principles into Bayesian inference for expensive models.
@@ 1.1 Background and motivation
We are specifically interested in exploring these methods for accelerating the calibration of computationally expensive climate models, because this type of calibration task exemplifies the challenges of combining UQ with real-world applications. Climate models play a crucial role in enhancing our understanding of climatological phenomena, allowing us to monitor current and past climatic conditions and forecast future trends [52, 25, 46]. Over time, these models have been refined to incorporate a multitude of intricate components, including wind patterns, atmospheric conditions, and more [51, 14, 30]. The utility of climate models, however, hinges on parameter settings that influence complex, interconnected processes across the atmosphere, oceans, and land surfaces. Calibrating physical and empirical parameters with UQ is critical for obtaining realistic model simulations and reducing climate model uncertainty. Calibration is especially challenging for climate models due to their complexity, computational demands, and the chaotic nature of the underlying systems, which amplifies uncertainty and hinders reliable parameter estimation.
Similar to other domains that depend on expensive physics models, calibration of climate models has historically been accomplished by manually setting parameters such that model output best matches observations, a task that requires the expertise of climate scientists [49]. This approach may introduce bias, ignore variability and uncertainty in model outputs [49], and fail to adequately capture parameter interactions. Recent research in climate calibration has sought to automate the calibration process. For instance, [59] proposes a scalable approach in which an emulator is trained on a limited number (250) of climate model simulations. The trained emulator is then used within an optimization procedure to determine optimal parameter values based on observations. The method proposed by [59] falls into a broad category of approaches called Bayesian calibration, in which parameter estimates and their uncertainty are rigorously quantified via a probability distribution, rather than a fixed value. Since the seminal paper that introduced the formalism of Bayesian calibration of computational models in 2001 [26], interest in this method has grown steadily in several fields [53], including biology [23], engineering [15, 17], physics [34] and climate science [24, 11, 6, 56, 27, 59, 50].
Efficient emulators are an important component of a Bayesian calibration framework, allowing it to scale to larger parameter spaces. In contrast to expensive models, emulators can be evaluated efficiently thousands to millions of times, which is necessary for Markov Chain Monte Carlo (MCMC) sampling typically involved in Bayesian calibration. A critical challenge in Bayesian calibration is selecting the limited set of model runs needed to train the emulator while maximizing its accuracy for calibration tasks. Traditional sampling strategies, such as Monte Carlo sampling and space-filling designs like Quasi Monte Carlo [42] and Latin hypercube sampling [38] improve efficiency compared to dense grids but are not specifically tailored to the downstream goal of calibration.
@@
Methods such as History Matching (HM) [57] and Calibrate, Emulate, Sample (CES) [7] aim to address this gap by focusing sampling on regions of the parameter space relevant to calibration. HM iteratively refines the parameter space by eliminating implausible regions, dramatically reducing computational costs [57]. It does not provide UQ on its own, although it can be combined with MCMC sampling to produce a posterior distribution over calibration parameters, which is the approach we take. CES incorporates information from the calibration problem into its sampling strategies. CES is unique in that it combines Kalman filtering, GP emulation, and MCMC sampling to iteratively refine the posterior, providing systematic UQ at each stage [7].
@@
Leveraging the formalism of Bayesian theory, BOED and GBOED extend these ideas by framing the design of model evaluations as an optimization problem. Within its standard formulation, BOED selects simulation designs that maximize the expected information gained about the emulator, leveraging the probabilistic nature of GPs to guide decisions. GBOED refines this approach further by focusing on regions of the parameter space that directly reduce uncertainty in the posterior distribution, making it particularly well-suited for calibration tasks. In essence, GBOED works by resolving intermediate calibration problems, leveraging the surrogate and its associated uncertainty. While BOED has been shown to achieve reliable calibration results for climate models with fewer model evaluations [12], the use of GBOED for calibration-specific tasks has not been explored in the literature.
@@
This paper is structured as follows: first, we review GP emulators (Section 2.1) and then provide an overview of the main methods (BOED, GBOED, CES, and HM) (Sections 2.2-2.4). Next, we describe the experimental setup (Section 3.2), validation (Section 3.1), and the model on which we test our methods (Section 4.2.1). Finally, we provide results regarding performance of all the methods (Section 4.2.2) and concluding remarks (Section 5).
@@ 2 Methods
In this study, we assume we have access to simulations from a climate model $\mathcal{G}$, which is typically very computationally expensive. Due to the computational intensity of $\mathcal{G}$ we seek to approximate it with a cheap emulator whenever we need to evaluate it many times, e.g. for model calibration. Typically the emulator does not capture the full output of $\mathcal{G}$, e.g. a very high dimensional spatiotemporal dataset, but only a reduced dimensional output comprised of important quantities of interest for solving the calibration problem. Since output quantities of interest from $\mathcal{G}$ are usually $p>1$ dimensional, two types of emulators may be considered. If outputs can be reasonably considered independent of one another, $p$ individual emulators can be used for modeling each component output. Alternatively, a multi-output emulator can be used to model correlation among the $p$ component outputs. Here, we consider independent GP emulators for each output dimension for reduced computational burden. Gaussian processes (GPs) are a natural choice for emulation due to their well-calibrated uncertainty quantification (UQ), flexibility, and analytical tractability, making them particularly well-suited for Bayesian calibration [55]. However, a wide variety of emulators have been used in the literature, including deep neural networks [3, 36], polynomial chaos expansions [39], and other machine learning methods [54].
@@ 2.1 Bayesian Calibration with GP emulators
Here, we provide a brief review of GPs for surrogate modeling. For more details see [55] or [22]. Consider the expensive computer model $\mathcal{G}$ represented as a function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{p}$ that maps inputs to outputs, whose pairs over $n$ simulations comprise data $D=(\mathbf{X},\mathcal{G}(\mathbf{X}))$, where inputs (model parameters) are $\mathbf{X}=\{x_{1}^{\top},x_{2}^{\top},\dots x_{m}^{\top}\}\in\mathbb{R}^{m\times d}$ and $\mathcal{G}(\mathbf{X})=\{\mathcal{G}(x_{1})^{\top},\mathcal{G}(x_{2})^{\top},\dots\penalty 0\mathcal{G}(x_{m})^{\top}\}\in\nobreak\mathbb{R}^{m\times p}$ are corresponding model outputs. Since we use independent GP emulators for each output dimension, we consider modeling the individual components $\mathcal{G}_{i}(\mathbf{X}),i=1\dots p$. For notational simplicity, let $\mathbf{Y}$ represent a single component output of $\mathcal{G}(\mathbf{X})$. Assuming a zero-mean GP prior for each component output of $f$ implies that $\mathbf{Y}$ follows a multivariate normal distribution (MVN) with a mean of zero and and a covariance $\Gamma_{GP}$ that is determined by inverse distances between inputs $\mathbf{X}$:
@@
The $\boldsymbol{\gamma}=(\gamma_{1},\gamma_{2},\dots\gamma_{d})$ are lengthscale or range parameters, where $1/\gamma_{\ell}$ determines the “relevance” of the $x_{\ell}$ input, $\tau^{2}$ governs the magnitude of the response, and the nugget $g$ divides the magnitude between signal $(\tau^{2})$ and noise $(\tau^{2}g)$. The GP model in Eq. (3) for $\mathbf{Y}$ gives rise to a MVN likelihood:
@@
$p(y_{obs}\mid\theta,\mathcal{GP}(D))=\frac{e^{-\frac{1}{2}(y_{obs}-\mu_{GP}(\theta))^{T}(\Gamma_{GP}(\theta)+\Gamma_{obs})^{-1}(y_{obs}-\mu_{GP}(\theta))}}{(2\pi)^{n/2}|\Gamma_{GP}(\theta)+\Gamma_{obs}|^{1/2}}.$ (5)
Because obtaining predictions from the GP emulator at any new inputs is cheap, MCMC can be used to perform Bayesian calibration via Eq. (5) while incorporating the model discrepancy due to the use of a GP emulator. The key challenge is judiciously choosing simulation training data $D$ to produce the most accurate and efficient emulator to be used in the calibration problem.
@@ 2.2 Bayesian Optimal Experimental Design
$EIG(\mathbf{X})$ $=\int p(\mathbf{Y}|\mathbf{X})\int p(\theta|y_{obs},\mathcal{GP}(D))\log\frac{p(\theta|y_{obs},\mathcal{GP}(D))}{p(\theta)}d\theta d\mathbf{Y}$
@@
In Eq. (2.2), $\mathbf{Y}|\mathbf{X}$ are realizations from the GP prior. In addition, $p(\theta|y_{obs},\mathcal{GP}(D))$ is the posterior distribution of parameters $\theta$, $p(\theta)$ is the prior distribution on $\theta$, and $p(y_{obs}|\theta,\mathcal{GP}(D))$ is the likelihood. We wish to choose the design $\mathbf{X}^{*}$ that maximizes the $EIG$; that is, given bounds on the input space ($\Xi$), we wish to find $\mathbf{X}^{*}=\arg\max_{\mathbf{X}\in\Xi}EIG(\mathbf{X})$. The expression in 2.2 represents the “idealized” OED problem and is rarely implemented in practice. Computing the EIG in the idealized OED problem is challenging, because for each candidate design $\mathbf{X}$, we must sample realizations $\mathbf{Y}$ from the GP prior predictive distribution, solve the calibration problem to obtain $p(\theta|y_{obs},\mathcal{GP}(D))$, and estimate the KL divergence or the marginal model evidence $p(y_{obs}|\mathcal{GP}(D))$.
To solve the idealized OED problem in Eq. (2.2), a sequential approach may be taken, where we iteratively optimize the EIG to determine an initial design $\mathbf{X}_{0}$, run simulations, and then conditionally optimize additional simulations. For the updated training dataset, we include new points $\mathbf{X}_{1}={\mathbf{X}_{0}\cup\mathbf{X}_{new}}$, where $\mathbf{X}_{new}$ represents newly selected inputs. This iterative process ensures that simulation samples are strategically placed to improve calibration performance of the emulator, rather than being driven by uninformed priors. Without this sequential refinement, samples are less effective, as they rely solely on prior information that may be uninformative. However, sequential refinement is computationally expensive, as it requires repeatedly solving the calibration problem and updating the emulator. In what follows, we emphasize two practical experimental design strategies: standard BOED, which maximizes EIG about the emulator, and goal-oriented BOED (GBOED), which prioritizes information gain about the inferred parameters. For completeness, we describe the “idealized” BOED formulation as a conceptual gold standard (though computationally infeasible for real-world problems) and compare it against BOED and GBOED in SM4.
@@
$EIG(\mathbf{X})=\int p(\mathbf{Y}|\mathbf{X})\int p(\mathbf{Y}^{\prime}|\mathbf{X^{\prime}},\mathcal{GP}(\mathbf{X},\mathbf{Y}))\log\frac{p(\mathbf{Y}^{\prime}|\mathbf{X^{\prime}},\mathcal{GP}(\mathbf{X},\mathbf{Y}))}{p(\mathbf{Y}^{\prime}|\mathbf{X}^{\prime})}\,dx\,d\mathbf{Y}.$ (7)
This expression quantifies how much we expect the GP posterior to differ from the GP prior at candidate locations $\mathbf{X}^{\prime}$, on average. Let the (zero-mean) GP prior have covariance matrix $\mathbf{K}_{X^{\prime}X^{\prime}}$, and let the posterior predictive variance at $\mathbf{X}^{\prime}$ after observing data $\mathcal{D}=(\mathbf{X},\mathbf{Y})$ be denoted by $\mathbf{\Sigma}^{\prime}$. Then, the EIG in (7) can be expressed in closed form as the expected KL divergence (over $\mathbf{Y}$) between two multivariate Gaussian distributions (Eq. 8) (see proof in SM5):
@@
In practice, not all GP outputs are equally relevant to calibration. To further improve efficiency, we introduce a goal-oriented BOED (GBOED) strategy, which focuses on predicting values most relevant to the calibration posterior. Instead of choosing a generic set of test inputs $\mathbf{X}^{\prime}$, GBOED utilizes test locations drawn from the current posterior of $\theta$, i.e., $\mathbf{X}^{\prime}\sim p(\theta|y_{obs},\mathcal{GP}(D))$. The GP is then optimized to gain information only about predictions at those points likely under the posterior, thereby aligning the training design more closely with the calibration task (see Algorithm 2.2). While BOED and GBOED differ in the choice of $\mathbf{X}^{\prime}$, we note that if $\mathbf{X}^{\prime}$ becomes dense in the input space (i.e., fully explores the emulator domain), then GBOED recovers standard BOED. We prove this in the finite case (see SM1). This theoretical equivalence justifies GBOED as a goal-oriented simplification of BOED that prioritizes the most relevant parts of the input space for calibration. For the interested reader, SM4 shows the differences in BOED, GBOED and idealized OED when applied to a toy calibration problem.
@@
Input: Bounds for input space $\Xi$, initial number of samples $n_{0}$, simulation budget $N$, GP emulator GP

### v2

@@ Abstract
Parameter calibration is essential for reducing uncertainty and improving predictive fidelity in physics-based models, yet it is often limited by the high computational cost of model evaluations. Bayesian calibration methods provide a principled framework for combining prior information with data while rigorously quantifying uncertainty. In this work, we compare four emulator-based Bayesian calibration strategies—Calibrate–Emulate–Sample (CES), History Matching (HM), Bayesian Optimal Experimental Design (BOED), and a goal-oriented extension of BOED (GBOED)—using the Lorenz ’96 multiscale system and a two-layer quasi-geostrophic (QG) model as test cases. The proposed GBOED formulation explicitly targets information gain with respect to the calibration posterior, aligning design decisions with downstream inference. We assess methods using accuracy and uncertainty quantification metrics, convergence behavior under increasing computational budgets, and practical considerations such as implementation complexity and robustness.
For the Lorenz ’96 system, CES, HM, and GBOED all yield strong calibration performance, even with limited numbers of model evaluations, while standard BOED generally underperforms in this setting. Differences among the strongest methods are modest, particularly as computational budgets increase. For the QG system, all methods produce reasonable posterior estimates, and convergence behavior is more consistent, reflecting smoother parameter–statistic relationships and improved identifiability. Overall, our results indicate that multiple emulator-based calibration strategies can perform comparably well when applied appropriately, with method selection often guided more by computational and practical considerations than by accuracy alone. These findings highlight both the limitations of standard BOED for calibration and the promise of goal-oriented and iterative approaches for efficient Bayesian inference in complex dynamical systems.
@@ 1 Introduction
Uncertainty quantification (UQ) is essential for validating predictive models in physics-based systems, where parameter choices propagate through nonlinear, often chaotic dynamics. Calibration, the process of inferring parameters from data with quantified uncertainty, is a many-query problem: it requires repeated evaluations of expensive simulators and thus demands methods that are both statistically reliable and computationally efficient.
A common strategy is to replace the simulator with an emulator, such as a Gaussian process (GP), that provides fast predictions with uncertainty. Emulator accuracy in calibration hinges on how simulation runs are chosen: informative design points can dramatically reduce cost, whereas uninformed sampling wastes evaluations and degrades posterior inferences.
We study Bayesian calibration for chaotic dynamical systems using emulators and compare four data-acquisition paradigms: Bayesian Optimal Experimental Design (BOED) [47, 9, 28] and its goal-oriented extension (GBOED) [3, 1, 78], Calibrate–Emulate–Sample (CES) [11], and History Matching (HM) [15]. While BOED has a long history in physical experimentation and sensor placement [75, 63, 26], its use for selecting computer simulations targeted to emulator-based calibration remains comparatively underexplored.
Contributions
We make the following contributions:
  1. 1.
    We introduce a goal-oriented BOED (GBOED) algorithm that directly targets information gain about the calibration posterior, rather than about the emulator alone, making the design criterion explicitly aligned with downstream inference.
  2. 2.
    We provide a comparative evaluation of BOED, GBOED, CES, and HM on the multiscale Lorenz’96 multiscale system, quantifying trade-offs among computational cost, implementation complexity, and posterior accuracy/uncertainty.
  3. 3.
    We assess the most promising methods on a computationally demanding quasi-geostrophic (QG) model, demonstrating performance and practical considerations in a higher-fidelity fluid-dynamics setting.
Results from applying the main methods to the two systems clarify when BOED and GBOED-based methods offer advantages for calibration, and how they compare to CES and HM when emulators drive the final MCMC-based inference.
@@ 1.1 Background and motivation
All methods considered in this study fall under the category of Bayesian calibration, in which parameter estimates and their uncertainty are rigorously quantified via a probability distribution, rather than a fixed value. From the Bayesian inverse-problems viewpoint, ensemble-based samplers provide optimization-free routes to posterior exploration. The ensemble Kalman sampler (EKS) interprets ensemble updates as a gradient flow for a mean-field approximation to the posterior [27], while ensemble Kalman inversion (EKI) adapts Ensemble Kalman Filter ideas to static inverse problems and has been analyzed/extended for nonlinear forward maps and regularization [37, 44, 67, 7, 36]. We are most interested in emulator-driven calibration, in which a surrogate (often a Gaussian process model) replaces an expensive model [42, 71, 32]. Emulators enable the thousands of forward evaluations required by uncertainty quantification (UQ) procedures such as Markov chain Monte Carlo (MCMC), while keeping total compute manageable. In some inference settings, emulators are trained in an amortized fashion, with the goal of supporting many downstream inference tasks across varying observational data [77]. However, calibration to a fixed observational data set does not fall into this category: accuracy is only required in regions of parameter space consistent with the observed data. This distinction motivates adaptive, calibration-aware design strategies that concentrate simulation effort where it is most informative for posterior inference, rather than seeking globally accurate surrogates. Since the seminal paper that introduced the formalism of Bayesian calibration of computational models in 2001 [42], interest in this method has grown steadily in several fields [69], including biology [33], engineering [21, 23], physics [48] and climate science [38, 19, 10, 73, 43, 76, 65].
Here, we compare the performance of four emulator-based calibration methods (CES, HM, BOED, GBOED) on calibrating two chaotic dynamical systems: the Lorenz ’96 multiscale system and the quasi-geostrophic (QG) system. We focus on systems where the forward model is moderately expensive and exhibits sensitive dependence on parameters and initial conditions, a combination that makes calibration challenging. We specifically study how different design and calibration choices interact in the regions of parameter space that matter most for inference.
@@
The design challenge in Bayesian calibration refers to selecting the limited set of model runs needed to train the emulator while maximizing its accuracy for calibration tasks. Traditional sampling strategies, such as Monte Carlo sampling and space-filling designs like Quasi Monte Carlo [56] and Latin hypercube sampling [52] improve efficiency compared to dense grids but are not specifically tailored to the downstream goal of calibration. Methods such as History Matching (HM) [74] and Calibrate, Emulate, Sample (CES) [11] aim to address this gap by focusing sampling on regions of the parameter space relevant to calibration. HM iteratively refines the parameter space by eliminating implausible regions, dramatically reducing computational costs [74]. HM does not provide UQ on its own, although it can be combined with MCMC sampling to produce a posterior distribution over calibration parameters, which is the approach we take. CES is unique in that it combines a Kalman filtering/sampling step, GP emulation, and MCMC sampling to iteratively refine the posterior, providing systematic UQ at each stage [11].
@@
Leveraging the formalism of Bayesian theory, BOED and GBOED extend these ideas by framing the design of model evaluations as an optimization problem. Within its standard formulation, BOED selects simulation designs that maximize the expected information gained about the emulator, leveraging the probabilistic nature of GPs to guide decisions. GBOED refines this approach further by focusing on regions of the parameter space that directly reduce uncertainty in the posterior distribution, making it particularly well-suited for calibration tasks. In essence, GBOED works by resolving intermediate calibration problems, leveraging the surrogate and its associated uncertainty. While BOED has been shown to achieve reliable calibration results for expensive models with fewer model evaluations [20], the use of GBOED for calibration-specific tasks has not been explored in the literature.
@@
This paper is structured as follows: first, we review GP emulators (Section 2.1) and provide an overview of the main methods (BOED, GBOED, CES, and HM) (Sections 2.2-2.4). Next, we describe the experimental setups for the Lorenz ’96 (Section 3.2), validation (Section 3.1), and results (Section 3.3). We describe the quasi-geostrophic system and results in Section 4. Finally, we provide a discussion and concluding remarks in Section 5.
@@ 2 Methods
In this study, we assume we have access to simulations from a model $\mathcal{G}$, which is typically computationally expensive. Due to the computational intensity of $\mathcal{G}$ we seek to approximate it with a cheap emulator whenever we need to evaluate it many times for the purpose of model calibration. Typically the emulator does not capture the full output of $\mathcal{G}$, e.g. a very high dimensional spatiotemporal dataset, but only a reduced dimensional output comprised of important quantities of interest for solving the calibration problem. Since output quantities of interest from $\mathcal{G}$ are usually $p>1$ dimensional, two types of emulators may be considered. If outputs can be reasonably considered independent of one another, $p$ individual emulators can be used for modeling each component output. Alternatively, a multi-output emulator can be used to model correlation among the $p$ component outputs. Here, we consider independent GP emulators for each output dimension for reduced computational burden (see SM5 for a description of the decorrelation procedure applied to outputs prior to GP fitting). Gaussian processes (GPs) are a natural choice for emulation due to their well-calibrated uncertainty quantification (UQ), flexibility, and analytical tractability, making them particularly well-suited for Bayesian calibration [72]. However, for larger dimensional problems, GP approximations may be required; scalable GP approximations include techniques such as compactly supported kernels [30, 41], fixed-rank kriging [16] inducing points [2, 59, 34] local divide-and-conquer [24, 31] and the Vecchia/nearest-neighbor approximation [17, 68, 40, 39, 5]. For larger problems, alternative emulators can be used, including deep neural networks [4, 50], polynomial chaos expansions [53], and other machine learning methods [70]. Additionally, in applications such as full climate fields, the number of raw outputs can be enormous, and fitting independent GPs to each component is neither computationally nor statistically sensible. One standard approach is to work in a reduced output space: compute a basis from an ensemble of simulator runs (using principal components analysis or other low-rank decompositions), retain $k$ coefficients that explain most variability, and emulate only those coefficients while treating the truncation residual as additional noise.
@@ 2.1 Bayesian Calibration with GP emulators
Here, we provide a brief review of GPs for surrogate modeling. For more details see [72] or [32]. Consider the expensive computer model $\mathcal{G}$ represented as a function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{p}$ that maps inputs to outputs, whose pairs over $n$ simulations comprise data $D=(\mathbf{X},\mathcal{G}(\mathbf{X}))$, where inputs (model parameters) are $\mathbf{X}=\{x_{1}^{\top},x_{2}^{\top},\dots x_{m}^{\top}\}\in\mathbb{R}^{m\times d}$ and $\mathcal{G}(\mathbf{X})=\{\mathcal{G}(x_{1})^{\top},\mathcal{G}(x_{2})^{\top},\dots\penalty 0\mathcal{G}(x_{m})^{\top}\}\in\penalty 10000\mathbb{R}^{m\times p}$ are corresponding model outputs. Since we use independent GP emulators for each output dimension, we consider modeling the individual components $\mathcal{G}_{i}(\mathbf{X}),i=1\dots p$. For notational simplicity, let $\mathbf{Y}$ represent a single component output of $\mathcal{G}(\mathbf{X})$. Assuming a zero-mean GP prior for each component output of $f$ implies that $\mathbf{Y}$ follows a multivariate normal distribution (MVN) with a mean of zero and and a covariance $\Gamma_{GP}$ that is determined by inverse distances between inputs $\mathbf{X}$:
@@
The $\boldsymbol{\gamma}=(\gamma_{1},\gamma_{2},\dots\gamma_{d})$ are lengthscale or range parameters, where $1/\gamma_{\ell}$ determines the “relevance” of the $x_{\ell}$ input, $\tau^{2}$ governs the magnitude of the response, and the nugget $g$ divides the magnitude between signal $(\tau^{2})$ and noise $(\tau^{2}g)$. In this study, we use the Matérn kernel with smoothness level fixed at 3.5. This implies translation-invariant covariance (dependence only on separation). Other non-stationary kernels with input-dependent length-scales could be used if stationarity is inappropriate for a given application, but in this study we assume stationarity. The GP model in Eq. (3) for $\mathbf{Y}$ gives rise to a MVN likelihood:
@@
$p(y_{obs}\mid\theta,\mathcal{GP}(D))=\frac{e^{-\frac{1}{2}(y_{obs}-\mu_{GP}(\theta))^{T}(\Gamma_{GP}(\theta)+\Gamma_{obs})^{-1}(y_{obs}-\mu_{GP}(\theta))}}{(2\pi)^{d/2}|\Gamma_{GP}(\theta)+\Gamma_{obs}|^{1/2}}.$ (5)
Because obtaining predictions from the GP emulator at any new inputs is cheap, MCMC can be used to perform Bayesian calibration via Eq. (5) while incorporating the model discrepancy due to the use of a GP emulator. For each of the methods considered in this paper, the final posterior for $\theta$ is computed via an MCMC algorithm using the likelihood in Eq. (5). The key challenge, however, in using GPs as part of a calibration method is judiciously choosing simulation training data $D$ to produce the most accurate and efficient emulator to be used in the calibration problem.
@@ 2.2 Bayesian Optimal Experimental Design
$\textup{EIG}(\mathbf{X})$ $=\int p(\mathbf{Y}|\mathbf{X})\int p(\theta|y_{obs},\mathcal{GP}(D))\log\frac{p(\theta|y_{obs},\mathcal{GP}(D))}{p(\theta)}d\theta d\mathbf{Y}$
@@
In Eq. (2.2), $\mathbf{Y}|\mathbf{X}$ are realizations from the GP prior. In addition, $p(\theta|y_{obs},\mathcal{GP}(D))$ is the posterior distribution of parameters $\theta$, $p(\theta)$ is the prior distribution on $\theta$, and $p(y_{obs}|\theta,\mathcal{GP}(D))$ is the likelihood. We wish to choose the design $\mathbf{X}^{*}$ that maximizes the EIG; that is, given bounds on the input space ($\Xi$), we wish to find $\mathbf{X}^{*}=\arg\max_{\mathbf{X}\in\Xi}\textup{EIG}(\mathbf{X})$. The expression in 2.2 represents the “idealized” OED problem and is rarely implemented in practice. Computing the EIG in the idealized OED problem is challenging, because for each candidate design $\mathbf{X}$, we must sample realizations $\mathbf{Y}$ from the GP prior predictive distribution, solve the calibration problem to obtain $p(\theta|y_{obs},\mathcal{GP}(D))$, and estimate the KL divergence or the marginal model evidence $p(y_{obs}|\mathcal{GP}(D))$.
To solve the idealized OED problem in Eq. (2.2), a sequential approach may be taken, where we iteratively optimize the EIG to determine an initial design $\mathbf{X}_{0}$, run simulations, and then conditionally optimize additional simulations. For the updated training dataset, we include new points $\mathbf{X}_{1}={\mathbf{X}_{0}\cup\mathbf{X}_{new}}$, where $\mathbf{X}_{new}$ represents a newly selected batch of inputs, which can be run as an ensemble in parallel. This iterative process ensures that simulation samples are strategically placed to improve calibration performance of the emulator, rather than being driven by uninformed priors. Without this sequential refinement, samples are less effective, as they rely solely on prior information that may be uninformative. However, sequential refinement is computationally expensive, as it requires repeatedly solving the calibration problem and updating the emulator. In what follows, we emphasize two practical experimental design strategies: standard BOED, which maximizes EIG about the emulator, and goal-oriented BOED (GBOED), which prioritizes information gain about the inferred parameters. For completeness, we describe the “idealized” BOED formulation as a conceptual gold standard (though computationally infeasible for real-world problems) and compare it against BOED and GBOED in SM3.
@@
$\textup{EIG}(\mathbf{X})=\int p(\mathbf{Y}|\mathbf{X})\int p(\mathbf{Y}^{\prime}|\mathbf{X^{\prime}},\mathcal{GP}(\mathbf{X},\mathbf{Y}))\log\frac{p(\mathbf{Y}^{\prime}|\mathbf{X^{\prime}},\mathcal{GP}(\mathbf{X},\mathbf{Y}))}{p(\mathbf{Y}^{\prime}|\mathbf{X}^{\prime})}\,d\mathbf{Y^{\prime}}\,d\mathbf{Y}.$ (7)
In Eq. 7, $\mathbf{X}^{\prime}$ is a fixed set of test locations; during each experiment optimization stage we fix $\mathbf{X}^{\prime}$ (hence we abbreviate to $\textup{EIG}(\mathbf{X})$ below). The inner integral marginalizes the random vector $\mathbf{Y}^{\prime}$, and the outer integral averages over possible training responses $\mathbf{Y}$ drawn from the GP prior predictive at $\mathbf{X}$. The expression quantifies how much we expect the GP posterior to differ from the GP prior at candidate locations $\mathbf{X}^{\prime}$, on average. Let the (zero-mean) GP prior have covariance matrix $\mathbf{K}_{X^{\prime}X^{\prime}}$, and let the posterior predictive variance at $\mathbf{X}^{\prime}$ after observing data $\mathcal{D}=(\mathbf{X},\mathbf{Y})$ be denoted by $\mathbf{\Sigma}^{\prime}$. Then, the EIG in (7) can be expressed in closed form as the expected KL divergence (over $\mathbf{Y}$) between two multivariate Gaussian distributions (Eq. 8) (see proof in SM4):
@@
Not all emulator predictions are equally relevant to calibration. Let $m(\theta)$ and $\Sigma_{total}(\theta)$ denote the emulator mean and total covariance (including $\Gamma_{obs}$). Locally, the calibration information carried by the outputs is governed by how $m(\theta)$ changes with $\theta$ relative to $\Sigma_{total}(\theta)$. Outputs that are weakly sensitive to $\theta$ or highly noise-dominated contribute little. Thus, choosing new simulations solely to reduce emulator error everywhere in $\theta$-space can be misaligned with calibration. To address this, we introduce a goal-oriented BOED (GBOED) strategy, which focuses on predicting values most relevant to the calibration posterior. Instead of choosing a generic set of test inputs $\mathbf{X}^{\prime}$, GBOED utilizes test locations drawn from the current approximated posterior of $\theta$, i.e., $\mathbf{X}^{\prime}\sim p(\theta|y_{obs},\mathcal{GP}(D))$, thereby aligning the training design more closely with the calibration task (see Algorithm 2.2). While BOED and GBOED differ in the choice of $\mathbf{X}^{\prime}$, we note that if $\mathbf{X}^{\prime}$ becomes dense in the input space (i.e., fully explores the emulator domain), then GBOED recovers standard BOED. We prove this in the finite case (see SM1). This theoretical equivalence justifies GBOED as a goal-oriented simplification of BOED that prioritizes the most relevant parts of the input space for calibration. In principle, emulator miscalibration could bias the posterior toward regions where the GP is less accurate. However, in practice, each $\mathbf{X}_{new}$ is appended to the training set and the GP is refit, which reduces predictive error locally. Throughout, “standard BOED” denotes the closed-form emulator-EIG in Eq. (2.7) evaluated at a fixed $\mathbf{X}^{\prime}$. GBOED uses the same objective but sets $\mathbf{X}^{\prime}\sim p(\theta|y_{obs},\mathcal{GP}(D))$ to target calibration, whereas Eq. (2.6) is the idealized parameter-EIG we include for context. For the interested reader, SM4 shows the differences in BOED, GBOED and idealized OED when applied to a toy calibration problem.
@@
Input: Bounds for input space $\Xi$, initial number of samples $n_{0}$, batch size $B$, total simulation budget $N$, GP emulator GP